# Supercoiling DNA with a free end

### v1

@@ 1 Introduction
Here we instead consider an open (linear) DNA filament with one free end, and use coarse-grained Brownian dynamics simulations [18, 19] to study the response of the duplex to the application of a constant torque at the other end. We ask whether the continuous injection of twist can still drive a transition to a plectonemic phase despite the fact that the free end can swivel. Differently from the buckling case, such transition would be inherently out-of-equilibrium. Our setup is related to those considered in [20, 21], but differs in several ways. In Ref. [20] twist is injected at one end via a constant angular velocity, and hydrodynamic interactions are included for studying plectoneme friction in short DNA chains. Ref. [21] studied the morphology of a twistable polymer constrained to wrap around a straight rod. Our study complements these previous ones for both the setup and the goals. In addition, we focus on the build-up, spatial distribution and steady-state profile of supercoiling (twist and writhe) along the chain.
This free boundary problem is much less studied than its fixed boundary or stretched end counterparts, but it could in principle be realised in single molecule experiments [22, 23]. It may also be relevant to model in vivo genomic segments that are far from topological barriers that hinder supercoiling propagation. Additionally, from a purely theoretical perspective, it provides a natural setting for studying transport and diffusion of supercoiling [24, 9, 5, 25], as well as the competition between twist injection and supercoil escape from the polymer ends.
@@
Our work is structured as follows. In Section 2, we present the twistable polymer model, along with details about its implementation in LAMMPS. In Section 3 we overview the phenomenology of the model in the limiting case of a perfectly linear polymer chain, which provides a baseline for twist transport. In Section 4, we present the main results for a initially equilibrated configurations of the model DNA duplex. Finally, we conclude by summarizing our findings and discussing possible refinements of the model and experimental validations of the predictions.
@@ 2 The polymer model
To describe the dynamics of double-stranded DNA, we employ a coarse-grained bead-chain model within the framework of Langevin dynamics, following the approach initially proposed by Chirico and Langowski [18]. Each bead $i$ has a diameter $\sigma=2.5$ nm, corresponding to approximately 7.3 base pairs (bp), and the chain consists of $N$ beads in total. Bead $i$ is characterized by its Cartesian position $\mathbf{r}_{i}$ and an orthonormal triad of unit vectors $\{\mathbf{f}_{i},\mathbf{v}_{i},\mathbf{u}_{i}\}$ (see Fig. 1(a)) defining the local body frame. The frame allows us to account for bending, through changes of $\mathbf{u}_{i}$ along the contour, and twist, through rotations around $\mathbf{u}_{i}$. To mimic the properties of a DNA duplex, the vector $\mathbf{u}_{i}$ is kept aligned with the local backbone direction by a suitable potential described below.
@@
where $I_{b}=m\sigma^{2}/10$ is the moment of inertia of a spherical bead, $\boldsymbol{\omega}_{i}$ is the spin angular velocity, and $\boldsymbol{\xi}_{i}(t)$ is a Gaussian stochastic torque with zero mean and covariance $\langle\xi_{i,a}(t)\xi_{j,b}(t^{\prime})\rangle=2\gamma_{\rm rot}k_{B}T\delta_{ij}\delta_{ab}\delta(t-t^{\prime})$. The rotational friction is set to $\gamma_{\rm rot}=s\gamma I/m$ with $s=10/3$, ensuring the correct rotational diffusivity for a sphere. The deterministic torque $\boldsymbol{\tau}_{U,i}$ derives from the potential $U$. An additional external torque $\tau_{a}\mathbf{u}_{i}$ is applied exclusively to the first bead ($i=0$) to drive twist propagation along the chain. Note that the end to which the torque is applied is not fixed but can move through space. From now on, we refer to the applied torque as an active torque.
@@
Torsional stiffness is modelled using Euler angles $\alpha_{i}$, $\beta_{i}$, $\gamma_{i}$ that describe the relative rotation between consecutive local frames [18, 19]:
@@
A phase angle $\phi_{i}$ is defined for each bead as $\phi_{i}=\sum_{j=0}^{i}(\alpha_{j}+\gamma_{j})$, so that $\alpha_{i}+\gamma_{i}=\phi_{i+1}-\phi_{i}$. The total twist stored in the chain is then ${\rm Tw}=\frac{1}{2\pi}\sum_{i=0}^{N-1}(\alpha_{i}+\gamma_{i})$.
@@
with $\mathbf{H}_{i}=\frac{\mathbf{u}_{i+1}\times\mathbf{u}_{i}\cos(\alpha_{i}+\gamma_{i})-\mathbf{f}_{i+1}\times\mathbf{f}_{i}-\mathbf{v}_{i+1}\times\mathbf{v}_{i}}{1+\mathbf{u}_{i+1}\cdot\mathbf{u}_{i}}$. All quantities are expressed in units of $m$, $\sigma$, and $k_{B}T$. The molecular dynamics time unit is $t_{\rm MD}=\sigma\sqrt{m/k_{B}T}$, and a timestep of $0.01\,t_{\rm MD}$ is used. As customary, the friction coefficient is set to $\gamma=2$ [31], yielding $\gamma_{\rm rot}=0.6$. Equations of motion are integrated using the velocity-Verlet algorithm in LAMMPS [32], with custom force terms implemented following Ref. [19].
@@
Figure 1: (a) Initial conformation at $t=0$ of a $2{\rm kbp}$ polymer. The inset shows an enlargement of the first bead index, with the orthonormal set $\{\mathbf{f}_{i},\mathbf{v}_{i},\mathbf{u}_{i}\}$ sketched for the first two beads. Highlighted is the active torque vector $\tau_{a}$, which is applied only on the bead $i=0$ in the direction of $\mathbf{u}_{0}$ and approximately parallel to the tangential vector $\mathbf{t}_{0}$. The circles highlight the direction of rotation caused by $\tau_{a}$. (b) Snapshots of $2{\rm kbp}$ long polymers subject to increasing torques $\tau_{a}=0,1,2,3,6$ (from bottom to top), and for increasing simulation times (from left to right). Chains are colored by chain index, see color bar in (a).
3 The linear polymer limit
As a preliminary step, we consider the simplified case of a perfectly linear chain: this idealized limit could be achieved, for instance, by imposing an infinitely large bending rigidity. In this case, all bond vectors $\mathbf{t}_{i}$ and local axes $\mathbf{u}_{i}$ remain strictly parallel, and the spin angular velocity $\boldsymbol{\omega}_{i}$ is collinear with $\mathbf{u}_{i}$. Consequently, the phase angle $\phi_{i}$ evolves as $\dot{\phi}_{i}={\omega}_{i}$. Under these conditions, the rotational equation of motion, Eq. (2), simplifies significantly. Averaging over stochastic noise, we obtain:
@@ 3 The linear polymer limit
$\sin(\phi_{i+1}-\phi_{i})=-\frac{\tau_{a}}{\kappa_{t}}\left(1-\frac{i}{N}\right).$ (12)
Thus, the stationary angular velocity increases linearly with the applied torque and decreases inversely with chain length. For positive $\tau_{a}$, a negative twist accumulates along the filament, with magnitude proportional to $\tau_{a}/\kappa_{t}$ and decaying linearly from the driven end ($i=0$) to the free end (i=N1i=N-1).
@@
with twist diffusivity $D_{\rm tw}=k_{B}Tl_{t}/(2\hat{\gamma}_{\rm rot})$, and $\delta^{\prime}(s)$ being the derivative of the Dirac delta function.
@@ 4 The three-dimensional polymer case
Having established the behavior in the limit of a linear chain, we now consider the case in which the DNA chain can adopt a fully three-dimensional configuration due to the combination of torque application and thermal fluctuations (Fig. 1). This unlocks additional relaxation channels – local bending and plectoneme formation – that dramatically reshape the propagation and dissipation of torsional stress along the filament. Hereafter, we systematically investigate the structural and dynamical response of the polymer by varying both the applied active torque magnitude and the DNA length.
Configurations at different torques. Fig. 1(b) shows configurations of polymers of equal length ($2{\rm kbp}$) under different torque values $\tau_{a}$ (low to high values from bottom to top rows) and at various times (different columns, time increasing from left to right). We start from the same initially equilibrated configuration in the absence of torque, and, starting from time $t=0$, apply the active torque at one end. As time progresses, we find two markedly different behaviors. For $\tau_{a}<\tau_{th}$, with $\tau_{th}\sim 2$ an active torque threshold, the chain remains in a swollen phase, with its morphology essentially unaffected by the imposed torsional stress. Instead, for $\tau_{a}\geq\tau_{th}$ the stress drives the formation of a plectoneme near the driven end, which gradually expands over time. Eventually, other plectonemes can form further from the driven end. Typically, plectonemes become more compact with increasing active torque, and at long times, we reach a steady state with one (or, less frequently, more than one) persisting plectoneme (see e.g., $t=1\cdot 10^{6}t_{MD}$).
@@
Figure 2: (a) Radius of gyration as a function of time, for different torques $\tau_{a}=0,1,2,3,6$, and fixed chain length $N=2\rm kbp$.(b) Radius of gyration, $R_{gyr}$, for various chain lengths and as a function of the active torque $\tau_{a}$. $R_{gyr}$ is scaled by $N^{\nu}$, with $\nu=0.588$ the Flory exponent.
@@
Spin angular velocity and global chain rotation. We now examine the profile of the average spin angular velocity projected along the local chain direction, $\boldsymbol{\omega}_{i}\cdot\mathbf{u}_{i}$, computed in the stationary regime, to assess how effectively the spin imparted at the driven end ($i=0$) propagates along a three-dimensional polymer chain.
Fig. 3(a) shows this profile for the shortest chain ($N=2\,\rm{kbp}$) and three different values of torque $\tau_{a}$. For any $\tau_{a}$, the spin velocity from the point of torque application ($i=0$) towards a finite plateau value at the free end ($i=N$). This behavior is consistent with previous observations in linear chains [20]. We observe a systematic increase in the overall spin velocity with increasing $\tau_{A}$. Additionally, the decay becomes progressively slower at higher torque.
@@
represented by the dashed line and the solid lines, respectively, where $\omega_{L}$ is the predicted spin angular velocity for a linear chain given by Eq. 11.
Beyond local spinning about its contour, the active torque also induces a global orbital rotation of the entire molecule. To characterize this collective motion, we computed the orbital angular velocity $\boldsymbol{\Omega}$ with respect to the polymer center of mass, through the following formula:
@@
where $\mathbf{I}$ is the moment-of-inertia tensor and $\boldsymbol{L}$ is the global angular momentum.
Fig. 3(d) shows the component of $\boldsymbol{\Omega}$ projected along $\mathbf{u}_{0}$, the direction along which the global rotation occurs. Global orbital rotation is apparent for $\tau_{a}\geq\tau_{th}$, with all curves collapsing onto a master curve when considering $\boldsymbol{\Omega}$ inversely proportional to $N^{2}$.
Figure 3: (a) Spin angular velocity profiles in the direction tangential to the chain, $\boldsymbol{\omega}_{i}\cdot\mathbf{u}_{i}$ as a function of the chain fraction $i/N$, for different torques $\tau_{a}=1,3,6$ and at fixed chain length $2\mathrm{kbp}$. (b) $\boldsymbol{\omega}_{i}\cdot\mathbf{u}_{i}$ as a function of $i/N$ for different chain length at fixed torque $\tau_{a}=6$. (c) $\boldsymbol{\omega}_{0}\cdot\mathbf{u}_{0}$ (triangles), and $\boldsymbol{\omega}_{N}\cdot\mathbf{u}_{N}$, diamonds, as a function of the applied torque for various chain lengths. The solid lines are the theoretical values $\omega_{L}=\frac{\tau_{a}}{N\gamma_{rot}}$, Eq. 11, computed for the different chain lengths. The dashed black line is $2.3\cdot 10^{-2}\tau_{a}$. (d) Orbital angular velocity $\boldsymbol{\Omega}$ of the chain projected on the direction of the applied torque $\mathbf{u}_{0}$, as a function of the applied torque for various chain lengths. $\boldsymbol{\Omega}$ has been scaled by the chain length squared $N^{2}$.
@@
Fig. 4(a)-(c) shows the global accumulation of twist and bend over time, for N=N=2 kbp and several $\tau_{a}$. Note that we plot $\rm Tw$}, as a positive active torque yields a negative accumulation of twist, as found for the linear chain. We also found that in the initial regime ${\rm Tw}\propto t^{0.5}$, compatible with a diffusive behavior for the twist propagation, which is expected for a linear chain from Eq. (15). For both $\rm Tw$} and $\beta$ curves reach saturation, but for twist it occurs much faster ($t\sim 10^{3}t_{MD}$) than for bending ($t\sim 10^{5}t_{MD}$); the latter timescale is compatible with the saturation time of $R_{gyr}$. As the dynamics of bending depends on plectoneme nucleation events [35], we do not expect a specific exponent to govern its growth law over time.
@@
Regarding twist accumulation, we find that for low active torque the profiles decay linearly to zero with the chain fraction $i/N$ (except for $\tau_{a}=0$, where it is always zero), which is again expected from Eq.12. Instead, at high enough torque, the profile develops a non-linear elbow close to the linear end, which deviates from the linear chain prediction (see comparison with dashed lines, corresponding to Eq.12 of a linear chain).
The bend profile starts again at a constant value at $\tau_{a}=0$ and develops with increasing $\tau_{a}$ to a maximum near the active torque application point $i=0$. As the torque increases, the peak becomes more prominent and shifts towards $i=0$.
@@
Figure 5: (a) Local twist angle $\alpha_{i}+\gamma_{i}$ and (b) local bending angle $\beta_{i}$ as a function of chain fraction $i/N$, at different active torques and fixed chain length $2{\rm kbp}$. The dashed lines in (a) represent, for each active torque value, the theoretical prediction for a linear chain, Eq. (12).
@@
Plectoneme formation and writhe. We now characterize the plectonemes, structures forming for $\tau_{a}>\tau_{th}$, by measuring the writhe, i.e. the number of self-crossings formed by the polymer. For a discrete polymer chain, the local writhe for a bead $i$ can be defined as [5]
${\rm Wr}_{i}=\frac{2}{4\pi}\sum_{k=i-l_{w}}^{i}\sum_{l=i}^{i+l_{w}}\frac{|(\mathbf{t}_{k,k+1}\times\mathbf{t}_{l,l+1})\cdot(\mathbf{r}_{k}-\mathbf{r}_{l})|}{|\mathbf{r}_{k}-\mathbf{r}_{l}|^{3}},$ (20)
@@
which, for each bead $i$, evaluates the $l_{w}$ preceding and $l_{w}$ following beads and counts the number of crossings within this window. In our analysis, we fixed $l_{w}=150$. To identify plectonemic regions, while avoiding noise, we first identify separate regions where $|{\rm Wr}|>1$, and then for each region we take the plectoneme borders as values where $|{\rm Wr}|=0.5$.

### v2

@@ 1 Introduction
Here we instead consider an open DNA filament with one free end, and use coarse-grained Brownian dynamics simulations [18, 19] to study the response of the duplex to the application of a constant torque at the other end. We ask whether the continuous injection of twist can still drive a transition to a plectonemic phase despite the fact that the free end can swivel. Different from the buckling case, such a transition would be inherently out of equilibrium. Our setup is related to those considered in [20, 21], but differs in several ways. In Ref. [20], twist is injected at one end via a constant angular velocity, and hydrodynamic interactions are included for studying plectoneme friction in short DNA chains. Ref. [21] studied the morphology of a twistable polymer constrained to wrap around a straight rod. Our study complements these previous ones for both the setup and the goals. We use Langevin dynamics to study relatively long twistable polymers, parametrized after DNA, that can freely fluctuate in three dimensions while a constant torque is applied at one of the ends. For simplicity, hydrodynamic interactions are neglected. In addition, we focus on the build-up, spatial distribution and steady-state profile of supercoiling (twist and writhe) along the chain.
This free boundary problem is much less studied than its fixed boundary or stretched end counterparts, but it could, in principle, be realised in single molecule experiments. In fact, torque can be applied either via RNA polymerases [22, 23], optical or magnetic tweezers [24, 25], or via electrosmotic flow [26, 27], while imaging of resulting plectonemes can be detected via fluorescence imaging [23]. It may also be relevant to model in vivo genomic segments that are far from topological barriers that hinder supercoiling propagation. Additionally, from a purely theoretical perspective, it provides a natural setting for studying transport and diffusion of supercoiling [28, 9, 5, 29], as well as the competition between twist injection and supercoil escape from the polymer ends. The latter process can supposedly occur via multiple modes, making the efficiency of supercoil release from the free end a non-trivial dynamical question. In the first mode, the free end is assumed to spin without resistance and thus efficiently dissipate injected twist [30]. However, this modality can be considered the sole one only if the chain remains approximately cylindrical, does not bend and has a sufficiently high torsional rigidity  [28]; otherwise, if writhe and plectonemes develop along the chain, one might expect injected twist to be dissipated either via solid-body rotation, or via plectoneme diffusion past the chain end[20], with a dynamics that is expected to be much slower compared to pure twist diffusion [5].
@@
Our work is structured as follows. In Section 2, we present the twistable polymer model, along with details about its implementation in LAMMPS. In Section 3, we overview the phenomenology of the model in the limiting case of a perfectly straight polymer chain, which provides a baseline for twist transport. In Section 4, we present the main results for a initially equilibrated configurations of the model DNA duplex. Finally, we conclude by summarizing our findings and discussing possible refinements of the model and experimental validations of the predictions.
@@ 2 The polymer model
To describe the dynamics of double-stranded DNA, we employ a coarse-grained bead-chain model within the framework of Langevin dynamics, following the approach initially proposed by Chirico and Langowski [18]. Each bead $i$ has a diameter $\sigma=2.5$ nm, corresponding to approximately 7.3 base pairs (bp), and the chain consists of $N$ beads in total, with the bead index $i$ running from 0 to N1N-1. Bead $i$ is characterized by its Cartesian position $\mathbf{r}_{i}$ and an orthonormal triad of unit vectors $\{\mathbf{f}_{i},\mathbf{v}_{i},\mathbf{u}_{i}\}$ (see Fig. 1(a)) defining the local body frame. The frame allows us to account for bending, through changes of $\mathbf{u}_{i}$ along the contour, and twist, through rotations around $\mathbf{u}_{i}$. To mimic the properties of a DNA duplex, the vector $\mathbf{u}_{i}$ is kept aligned with the local backbone direction by a suitable potential described below.
@@
where $I_{b}=m\sigma^{2}/10$ is the moment of inertia of a spherical bead, $\boldsymbol{\omega}_{i}$ is the spin angular velocity, and $\boldsymbol{\xi}_{i}(t)$ is a Gaussian stochastic torque with zero mean and covariance $\langle\xi_{i,a}(t)\xi_{j,b}(t^{\prime})\rangle=2\gamma_{\rm rot}k_{B}T\delta_{ij}\delta_{ab}\delta(t-t^{\prime})$. The rotational friction is set to $\gamma_{\rm rot}=s\gamma I_{b}/m$ with $s=10/3$, ensuring the correct rotational diffusivity for a sphere. The deterministic torque $\boldsymbol{\tau}_{U,i}$ derives from the potential $U$. An additional external torque of fixed magnitude $\tau_{a}$ is applied exclusively to the first bead (i= 0) along its instantaneous axis $\mathbf{u}_{0}$ to drive twist propagation along the chain. Note that both chain ends, including the one to which the torque is applied are not clamped but can move through space. From now on, we refer to the applied torque as an active torque.
@@
Torsional stiffness is modeled using Euler angles $\alpha_{i}$, $\beta_{i}$, $\gamma_{i}$ that describe the relative rotation between consecutive local frames [18, 19], see Supplementary Figure S1:
@@
A phase angle $\phi_{i}$ is defined for each bead as $\phi_{i}=\sum_{j=0}^{i}(\alpha_{j}+\gamma_{j})$, so that $\alpha_{i}+\gamma_{i}=\phi_{i+1}-\phi_{i}$. The total twist stored in the chain is then ${\rm Tw}=\frac{1}{2\pi}\sum_{i=0}^{N-2}(\alpha_{i}+\gamma_{i})$.
@@
with $\mathbf{H}_{i}=\frac{\mathbf{u}_{i+1}\times\mathbf{u}_{i}\cos(\alpha_{i}+\gamma_{i})-\mathbf{f}_{i+1}\times\mathbf{f}_{i}-\mathbf{v}_{i+1}\times\mathbf{v}_{i}}{1+\mathbf{u}_{i+1}\cdot\mathbf{u}_{i}}$.
Note that in this model, the typical parameters to tune in order to reproduce the DNA properties (e.g., in specific ionic conditions) are the elasticity parameters, $l_{b}$ and $l_{t}$, and the effective filament thickness, $\sigma$. Here, such values are chosen to match DNA in physiological conditions, $\sigma=2.5$nm, $l_{b}=50nm,$ and $l_{t}=150nm$. All quantities are expressed in units of $m$, $\sigma$, and $k_{B}T$. The molecular dynamics time unit is $t_{\rm MD}=\sigma\sqrt{m/k_{B}T}$, and a timestep of $0.01\,t_{\rm MD}$ is used. As customary, the friction coefficient is set to $\gamma=2$ [35], yielding $\gamma_{\rm rot}=0.66$. Equations of motion are integrated using the velocity-Verlet algorithm in LAMMPS [36], with custom force terms implemented following Ref. [19].
@@
A nominal mapping from simulation time to real physical units can be performed as follows. The relevant microscopic characteristic timescale is given by the nominal self-diffusion time of a single monomer in the chain. Equivalently, one can define the mapping by matching the self-diffusion coefficient of an isolated bead $D=\frac{k_{B}T}{\gamma}=0.5\frac{\sigma^{2}}{t_{MD}}$ with the one estimated by applying Stokes’ law to a bead of diameter $\sigma=2.5nm$ immersed in a fluid of viscosity equal to water, 1cP. Setting $T=300K$, the matching, which neglects hydrodynamic interactions, yields $t_{MD}\approx 40$ ns. Thus, our simulations, which last typically $10^{8}t_{MD}$, correspond to $4$s.
Figure 1: (a) Initial conformation at $t=0$ of a $2{\rm kbp}$ polymer. The inset shows an enlargement of the first bead index, with the orthonormal set $\{\mathbf{f}_{i},\mathbf{v}_{i},\mathbf{u}_{i}\}$ sketched for the first two beads. Highlighted is the active torque vector $\tau_{a}$, which is applied only on the bead $i=0$ in the direction of $\mathbf{u}_{0}$ and approximately parallel to the tangential vector $\mathbf{t}_{0}$. The circles highlight the direction of rotation caused by $\tau_{a}$. (b) Snapshots of $2{\rm kbp}$ long polymers subject to increasing torques $\tau_{a}=0,1,2,3,6$ (from bottom to top), and for increasing simulation times (from left to right). Chains are colored by chain index, see color bar in (a). Supplementary Movies S1-S5 display the complete trajectories for the cases illustrated in (b).
3 Perfectly straight chain limit
As a preliminary step, we consider the simplified case of a perfectly straight chain: this idealized limit could be achieved, for instance, by imposing an infinitely large bending rigidity. In this case, all bond vectors $\mathbf{t}_{i}$ and local axes $\mathbf{u}_{i}$ remain strictly parallel, and the spin angular velocity $\boldsymbol{\omega}_{i}$ is collinear with $\mathbf{u}_{i}$. Consequently, the phase angle $\phi_{i}$ evolves as $\dot{\phi}_{i}={\omega}_{i}$. Under these conditions, the rotational equation of motion, Eq. (2), simplifies significantly. Averaging over stochastic noise, we obtain:
@@ 3 Perfectly straight chain limit
$\sin(\phi_{i}-\phi_{i-1})=-\frac{\tau_{a}}{\kappa_{t}}\left(1-\frac{i}{N}\right).$ (12)
Thus, the stationary angular velocity increases linearly with the applied torque and decreases inversely with chain length. For positive $\tau_{a}$, a negative twist accumulates along the filament, with magnitude proportional to $\tau_{a}/\kappa_{t}$. Note that the twist profile decays linearly from the driven end ($i=0$) to the free end (i=N1i=N-1) for $|\tau_{a}/\kappa_{t}|\leq 1$ (e.g. in a small-torque regime), where $\sin(\phi_{i+1}-\phi_{i})\approx\phi_{i+1}-\phi_{i}$.
@@
with twist diffusivity $D_{\rm tw}=k_{B}Tl_{t}/(2\hat{\gamma}_{\rm rot})$, and $\delta^{\prime}(s)$ being the derivative of the Dirac delta function. Note that for a finite polymer with boundaries $s\in[0,N\sigma]$, one needs to set zero accumulation of twist outside these boundaries, $\rho_{\rm tw}=0$. Equivalently, one can recast Eqs. 13 and  15 to their homogeneous forms (i.e. without involving the delta function) supplemented with the boundary conditions $\rho_{\rm tw}(0,t)=-\frac{\tau_{a}}{2\pi D_{\rm tw}\hat{\gamma}{\rm rot}}$, and $\rho_{\rm tw}(N\sigma,t)=0$.
@@ 4 The three-dimensional polymer case
Having established the behavior in the limit of a straight chain, we now consider the case in which the DNA chain can adopt a fully three-dimensional configuration due to the combination of torque application and thermal fluctuations (Fig. 1). This unlocks additional relaxation channels – local bending and plectoneme formation – that dramatically reshape the propagation and dissipation of torsional stress along the filament. Hereafter, we systematically investigate the structural and dynamical response of the polymer by varying both the applied active torque magnitude and the DNA length.
Configurations at different torques. Fig. 1(b) shows configurations of polymers of equal length ($2{\rm kbp}$) under different torque values $\tau_{a}$ (low to high values from bottom to top rows) and at various times (different columns, time increasing from left to right), see also Supplementary Movies S1-S5. We start from the same initially equilibrated configuration in the absence of torque, and, starting from time $t=0$, apply the active torque at one end. As time progresses, we find two markedly different behaviors. For $\tau_{a}<\tau_{th}$, with $\tau_{th}\sim 2$ an active torque threshold, the chain remains in a swollen phase, with its morphology essentially unaffected by the imposed torsional stress. Instead, for $\tau_{a}\geq\tau_{th}$ the stress drives the formation of a plectoneme near the driven end, which gradually expands over time. Eventually, other plectonemes can form further from the driven end. Typically, plectonemes become more compact with increasing active torque, and at long times, we reach a steady state with one (or, less frequently, more than one) persisting plectoneme (see e.g., $t=1\cdot 10^{6}t_{MD}$).
@@
Figure 2: (a) Radius of gyration as a function of time, for different torques $\tau_{a}=0,1,2,3,6$, and fixed chain length $N=2\rm kbp$.(b) Rescaled gyration radius, $R_{gyr}/N^{\nu}$, for various chain lengths and as a function of the active torque $\tau_{a}$. The Flory scaling exponent was set to $\nu=0.588$.
@@
Spin angular velocity and global chain rotation. We now examine the profile of the average spin angular velocity projected along the local chain direction, $\boldsymbol{\omega}_{i}\cdot\mathbf{u}_{i}$, computed in the stationary regime, to assess how effectively the spin imparted at the driven end ($i=0$) propagates along a three-dimensional polymer chain. In other words, it quantifies the extent to which the chain is rotating (spinning) about its own local tangent axis—i.e., torsional propagation along the backbone—see also Supplementary Figure S1.
Fig. 3(a) shows this profile for the shortest chain ($N=2\,\rm{kbp}$) and three different values of torque $\tau_{a}$. For any $\tau_{a}$, the spin velocity from the point of torque application ($i=0$) towards a finite plateau value at the free end ($i=N$). This behavior is consistent with previous observations in Ref. [20]. We observe a systematic increase in the overall spin velocity with increasing $\tau_{a}$. Additionally, the decay becomes progressively slower at higher torque.
@@
represented by the dotted line and the solid lines, respectively, where $\omega_{L}$ is the predicted spin angular velocity for a straight chain given by Eq. 11. Interestingly, in Ref. [20] the presence of hydrodynamics affects differently the scaling of $\boldsymbol{\omega}_{N}\cdot\mathbf{u}_{N}$, which remains equal to $\boldsymbol{\omega}_{0}\cdot\mathbf{u}_{0}$ in the swollen phase, and becomes proportional to $N^{-2/3}$ in the plectonemic phase.
Beyond local spinning about its contour, the active torque also induces a global orbital rotation of the entire molecule, where the entire coil whirls as a single unit around a rotational axis, see Supplementary Figure S1. To characterize this collective motion, we computed the orbital angular velocity $\boldsymbol{\Omega}$ with respect to the polymer center of mass, through the following formula:
@@
where $\mathbf{I}$ is the moment-of-inertia tensor and $\boldsymbol{L}$ is the global angular momentum. Note that we obtain $\boldsymbol{\Omega}$ by computing directly both $\mathbf{I}$ and $\boldsymbol{L}$ in a given conformation, and inverting the tensor relation.
Fig. 3(d) shows the component of $\boldsymbol{\Omega}$ projected along $\mathbf{u}_{0}$, the direction along which the global rotation occurs, and averaged at steady state over different conformations. Global orbital rotation is apparent for $\tau_{a}\geq\tau_{th}$, with all curves collapsing onto a master curve when rescaling $\boldsymbol{\Omega}$ by $N^{2}$, implying that $\boldsymbol{\Omega}\propto N^{-2}$ at fixed $\tau_{a}$.
Figure 3: (a) Spin angular velocity profiles in the direction tangential to the chain, $\boldsymbol{\omega}_{i}\cdot\mathbf{u}_{i}$ as a function of the chain fraction $i/N$, for different torques $\tau_{a}=1,3,6$ and at fixed chain length $2\mathrm{kbp}$. (b) $\boldsymbol{\omega}_{i}\cdot\mathbf{u}_{i}$ as a function of $i/N$ for different chain length at fixed torque $\tau_{a}=6$. (c) $\boldsymbol{\omega}_{0}\cdot\mathbf{u}_{0}$ (triangles), and $\boldsymbol{\omega}_{N}\cdot\mathbf{u}_{N}$, diamonds, as a function of the applied torque for various chain lengths. The solid lines are the theoretical values $\omega_{L}=\frac{\tau_{a}}{N\gamma_{rot}}$, Eq. 11, computed for the different chain lengths. The dotted black line, obtained from a fit, is $2.3\cdot 10^{-2}\tau_{a}$. (d) Orbital angular velocity $\boldsymbol{\Omega}$ of the chain projected on the direction of the applied torque $\mathbf{u}_{0}$, as a function of the applied torque for various chain lengths. $\boldsymbol{\Omega}$ has been scaled by the chain length squared $N^{2}$.
@@
Fig. 4(a)-(c) shows the global accumulation of twist and bend over time, for N=N=2 kbp and several $\tau_{a}$. Note that we plot $\rm Tw$}, as a positive active torque yields a negative accumulation of twist, as found for the straight chain. We also found that in the initial regime ${\rm Tw}\propto t^{0.5}$, compatible with a diffusive behavior for the twist propagation, which is expected for a straight chain from Eq. (15). For both $\rm Tw$} and $\beta$ curves reach saturation, but for twist it occurs much faster ($t\sim 10^{3}t_{MD}$) than for bending ($t\sim 10^{5}t_{MD}$); the latter timescale is compatible with the saturation time of $R_{gyr}$. Using the nominal time-to-physical-unit mapping, the indicative characteristic saturation times of twist and bending lie in the range 0.0440.04-4 ms, in agreement with Ref. [38]. As the dynamics of bending depends on plectoneme nucleation events [40], we do not expect a specific exponent to govern its growth law over time.
@@
Regarding twist accumulation, we find that for low active torque the profiles decay linearly to zero with the chain fraction $i/N$ (except for $\tau_{a}=0$, where it is always zero), which is again expected from Eq.12. Instead, at high enough torque, the profile develops a non-linear elbow close to the linear end, which deviates from the straight chain prediction (see comparison with dashed lines, corresponding to Eq.12 of a straight chain).
The bend profile starts again at a constant value at $\tau_{a}=0$ and develops with increasing $\tau_{a}$ to a maximum near the active torque application point $i=0$. As the torque increases, the peak becomes more prominent and shifts towards $i=0$. In a plectonemic conformation, which is inherently more rigid than the swollen state, the bending angle exceeds that arising from thermal fluctuations alone due to the approximately helical local structure of the plectoneme; in fact, although the overall plectoneme may appear relatively straight along its centerline, each of the two intertwined duplexes exhibits substantial curvature around this centerline, as each can be locally modeled as a helix with a radius of curvature determined by the pitch and projected radius of the plectoneme (see Supplementary Figure S2).
@@
Figure 5: (a) Local twist angle $\alpha_{i}+\gamma_{i}$ and (b) local bending angle $\beta_{i}$ as a function of chain fraction $i/N$, at different active torques and fixed chain length $2{\rm kbp}$. The dashed lines in (a) represent, for each active torque value, the theoretical prediction for a straight chain, Eq. (12).
@@
Plectoneme formation and writhe. We now characterize the plectonemes, structures forming for $\tau_{a}>\tau_{th}$, by measuring the writhe, i.e., the number of self-crossings formed by the polymer. For a discrete polymer chain, the local writhe for a bead $i$ can be defined as [5]
${\rm Wr}_{i}=\frac{2}{4\pi}\sum_{k=i-l_{w}}^{i}\sum_{l=i}^{i+l_{w}}\frac{(\mathbf{t}_{k,k+1}\times\mathbf{t}_{l,l+1})\cdot(\mathbf{r}_{k}-\mathbf{r}_{l})}{|\mathbf{r}_{k}-\mathbf{r}_{l}|^{3}},$ (20)
@@
which, for each bead $i$, evaluates the $l_{w}$ preceding and $l_{w}$ following beads and counts the number of crossings within this window. In our analysis, we fixed $l_{w}=150$. To identify plectonemic regions, while avoiding noise, we first identify separate regions where $|{\rm Wr_{i}}|>1$, and then extend them to smaller and larger values of $i$ until $|{\rm Wr_{i}}|$ falls below the threshold value $0.5$, thus defining the two boundaries. We can also define the total writhe as: [18]
${\rm Wr}=\frac{1}{4\pi}\sum_{k=0}^{N-2}\sum_{l=0}^{N-2}\frac{\left(\mathbf{t}_{k,k+1}\times\mathbf{t}_{l,l+1}\right)\cdot\left(\mathbf{r}_{k}-\mathbf{r}_{l}\right)}{|\mathbf{r}_{k}-\mathbf{r}_{l}|^{3}}.$ (21)